Suppose that for a very large shipment of​ integrated-circuit chips, the probability of failure for any one chip is 0.10.Assuming that the assumptions underlying the binomial distributions are​ met, find the probability that at most 2 chips fail in a random sample of 19.
Answer #1

Solution

Given that ,

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X 2) = P(x = 0) + P(x = 1) + P(x = 2)

= (19 C 0) * 0.100 * (0.90)19 + (19 C 1) * 0.101 * (0.90)18+ (19 C 2) * 0.102 * (0.90)17

= 0.7054

Probability = 0.7054